Solve the following equations.  7x – 9 = 16 
Solve the following equations.  7x – 9 = 16 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

7x–9=16
7x = 16+9
7x=25
x=25/7

Thanks